OpenStudy (anonymous):

sin3 x /cos x dx... can help me solve the exercise?

OpenStudy (zarkon):

expand sin(3x)

OpenStudy (anonymous):

I was wrong to write the equation , it really is sin^3 x /cos x dx

OpenStudy (zarkon):

use... \[\sin^3(x)=\sin(x)\sin^2(x)=\sin(x)(1-\cos^2(x))\] then use \(u\) substitution.